# The local homological structure of generalized splines

### v1

@@ 1 Introduction
Historically, this mirrors the early work in classical splines initiated by Billera and Rose, who first used a local–global argument for freeness based on stars [2, Theorem 2.3]. Later, Schenck and Stillman advanced this approach with the use of local cohomology and depth estimates to analyze spline modules [5]. More recently, Altınok and Sarıoğlan [1] developed a homological approach with generalized splines. The present work deviates from the classical heritage in two central ways. First, the entire situation is understood from the perspective of generalized splines, and second, generalized splines are also interpreted scheme-theoretically [7]. As a result, the local–global machinery and local cohomology are distilled so that the roles of algebra and combinatorics are clarified.
To briefly summarize the main results: there are two exact sequences that control the homological behavior of generalized splines over local rings. The first is the potential exact sequence, for which the splines are the kernel and the cosplines are the cokernel. The second is the copotential exact sequence, which has the cosplines as the kernel, and the failed cosplines as the cokernel. The objects of these two exact sequences are closely related to the vertex and edge structure of the associated edge-labeled graph, and thereby provide a convenient control point for proofs. To exemplify how to appropriately use the exact sequences, we prove a general theorem on Cohen–Macaulay rings with certain properties. As a special case, the theorem shows that any multivariate generalized spline module in two variables with principal edge labels is free.
@@ 2 The Basic Homology of Local Rings
$\operatorname{pd}_{R}(M)+\operatorname{depth}_{R}(M)=\operatorname{depth}(R).$
@@
where $\operatorname{Ext}$ is the derived functor cohomology of Hom [6, Lemma 10.72.5]. We can also work equivalently by local cohomology:
@@
where $H$ is the derived functor cohomology of the $\mathfrak{m}$-torsion functor [6, Lemma 47.11.1]. Furthermore, for finite direct sums we have
@@
provided the $M_{i}$ are finitely generated modules [4, Example 1.8]. For Cohen–Macaulay rings $R$, the Krull dimension of $R$ and the depth of $R$ as an $R$-module over itself coincide, which allows the $\operatorname{depth}(R)$ portion of the Auslander–Buchsbaum formula to be replaced with $\dim(R)$. Critically, for local Noetherian Cohen–Macaulay rings, the depth drop formula has $\operatorname{depth}R/\langle f\rangle=\operatorname{depth}R-1$ for $\mathfrak{m}$-regular $f\in R$, where the depth of the quotient ring is as an $R$-module [4, Proposition 2.9]. As we will see later, this minor and predictable depth drop is precisely why principal ideal labels are favorable to the presence of free generalized spline modules.
@@
We also know that an $R$-module $M$ (again, $R$ is local) has depth 0 if and only if $\mathfrak{m}$ is an associated prime of $M$, that is, every element of $\mathfrak{m}$ is a zero divisor on $M$ [4, Example 1.2]. Furthermore, we have that the short exact sequence:
@@
per [6, Lemma 10.63.3]. We thus have two mechanisms of depth control for $N^{\prime}$ via $N$ and $N^{\prime\prime}$. The first is to provide a lower bound for the depth of $N^{\prime}$. The second is to observe that if $N$ has depth $1$, then $\mathfrak{m}$ is not an associated prime of $N$, and since $N^{\prime}$ is a submodule, $\mathfrak{m}$ cannot be an associated prime of $N^{\prime}$. This means $N^{\prime}$ cannot have depth 0. This allows $N$ to regulate the depth of $N^{\prime}$ directly, albeit in a fairly restricted way. The other mechanism requires both $N$ and $N^{\prime\prime}$ working together, but provides a more robust floor on the depth of $N^{\prime}$.
@@ 3 The Homological Theory of Generalized Splines
In particular, when $R$ is Noetherian, the finiteness of $G$ implies that the modules $R^{V}$, $R^{E}$, $R_{G}$, $\overline{R_{G}}$, and $\underline{R_{G}}$ are finitely generated $R$-modules. These give rise to two fundamental exact sequences: the potential exact sequence and copotential exact sequence.
@@ 4 Applications
Since $R$ is a domain, every non-zero principal ideal is generated by a regular element. Let $\mathfrak{m}$ be a maximal ideal of $R$. Now by the stability of spline structure morphisms under flat base change [7, Main Theorem] we have that $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$. That is, $(R_{G})_{\mathfrak{m}}$ is isomorphic to a module of splines over the local ring $R_{\mathfrak{m}}$ that has edge-labeled graph $G_{\mathfrak{m}}$. By hypothesis $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ satisfies the conditions of Theorem 3.5, and is therefore free.
Furthermore, each $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ is the kernel of a potential short exact sequence, for which the middle term is free, and since each $R_{\mathfrak{m}}$ is Noetherian [6, Lemma 10.31.1] we have that $(R_{G})_{\mathfrak{m}}$ is finitely generated. Now for modules over Noetherian rings, finitely generated implies finitely presented [6, Lemma 10.31.4] and therefore each $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ is also finitely presented.
So we have established that for each maximal ideal $\mathfrak{m}$, we have that $(R_{G})_{\mathfrak{m}}$ is free and of finite presentation. Now any $R$-module that is free and of finite presentation locally at each maximal ideal of $R$ is also finite projective [6, Lemma 10.78.2(5)], so it follows that $R_{G}$ is a finite projective $R$-module as claimed. ∎
@@
Let $R$, $G$, and $R_{G}$ be given as specified. Note that by hypothesis $R$ is the coordinate ring of a nonsingular surface, and is therefore a regular ring [6, Definition 28.9.1, Lemma 28.9.2(3)]. Since regular rings are Cohen–Macaulay [4, Example 2.2], we can see $R$ is Cohen–Macaulay. We also know that $R$ is Krull dimension $2$ at each maximal ideal by hypothesis. Furthermore, since $R$ is a regular ring, it has finite global dimension, which implies that $R_{G}$ has finite projective dimension [6, Lemma 10.110.8, Definition 10.109.2]. Finally, by the irreducibility hypothesis we know $R$ is an integral domain, and it is Noetherian since the quotient of a Noetherian ring is Noetherian. So the criteria of Theorem 4.1 are satisfied and the claim follows. ∎

### v2

@@ 1 Introduction
Historically, this mirrors the early work in classical splines initiated by Billera and Rose, who first used a local–global argument for freeness based on stars [2, Theorem 2.3]. Later, Schenck and Stillman advanced this approach with the use of local cohomology and depth estimates to analyze spline modules [6]. In the field of generalized splines, Altınok and Sarıoğlan [1] developed a homological approach to multivariate generalized splines using syzygies, while Michael DiPasquale set up a homological framework for generalized splines using an exact sequence [4]. The approach taken here follows the homological direction already initiated for generalized splines, but with a focus on local rings. This is motivated by the interpretation of generalized splines as schemes, which facilitates a local–global principle intrinsic to the ring itself [8, Main Theorem].
To briefly summarize the main results: there are two exact sequences that control the homological behavior of generalized splines over local rings. The first is the potential exact sequence, for which the splines are the kernel and the cosplines are the cokernel, which was first studied by DiPasquale [4, Lemma 2.2]. The second is the copotential exact sequence, which has the cosplines as the kernel, and the failed cosplines as the cokernel. The objects of these two exact sequences are closely related to the vertex and edge structure of the associated edge-labeled graph, and thereby provide a convenient control point for proofs. To exemplify how to appropriately use the exact sequences, we prove a general theorem on Cohen–Macaulay rings with certain properties. As a special case, the theorem shows that any multivariate generalized spline module in two variables with principal edge labels is free.
@@ 2 The Basic Homology of Local Rings
$\operatorname{pd}_{R}(M)+\operatorname{depth}_{R}(M)=\operatorname{depth}_{R}(R).$
@@
where $\operatorname{Ext}$ is the derived functor cohomology of Hom [7, Lemma 10.72.5]. We can also work equivalently by local cohomology:
@@
where $H$ is the derived functor cohomology of the $\mathfrak{m}$-torsion functor [7, Lemma 47.11.1]. Furthermore, for finite direct sums we have
@@
provided the $M_{i}$ are finitely generated modules [5, Example 1.8]. For Cohen–Macaulay rings $R$, the Krull dimension of $R$ and the depth of $R$ as an $R$-module over itself coincide, which allows the $\operatorname{depth}(R)$ portion of the Auslander–Buchsbaum formula to be replaced with $\dim(R)$. Critically, for local Noetherian Cohen–Macaulay rings, the depth drop formula has $\operatorname{depth}R/\langle f\rangle=\operatorname{depth}R-1$ for $\mathfrak{m}$-regular $f\in R$, where the depth of the quotient ring is as an $R$-module [5, Proposition 2.9]. As we will see later, this minor and predictable depth drop is precisely why principal ideal labels are favorable to the presence of free generalized spline modules.
@@
We also know that an $R$-module $M$ (again, $R$ is local) has depth 0 if and only if $\mathfrak{m}$ is an associated prime of $M$, that is, every element of $\mathfrak{m}$ is a zero divisor on $M$ [5, Example 1.2]. Furthermore, we have that the short exact sequence:
@@
per [7, Lemma 10.63.3]. We thus have two mechanisms of depth control for $N^{\prime}$ via $N$ and $N^{\prime\prime}$. The first is to provide a lower bound for the depth of $N^{\prime}$. The second is to observe that if $N$ has depth $1$, then $\mathfrak{m}$ is not an associated prime of $N$, and since $N^{\prime}$ is a submodule, $\mathfrak{m}$ cannot be an associated prime of $N^{\prime}$. This means $N^{\prime}$ cannot have depth 0. This allows $N$ to regulate the depth of $N^{\prime}$ directly, albeit in a fairly restricted way. The other mechanism requires both $N$ and $N^{\prime\prime}$ working together, but provides a more robust floor on the depth of $N^{\prime}$.
@@ 3 The Homological Theory of Generalized Splines
In particular, when $R$ is Noetherian, the finiteness of $G$ implies that the modules $R^{V}$, $R^{E}$, $R_{G}$, $\overline{R_{G}}$, and $\underline{R_{G}}$ are finitely generated $R$-modules. These give rise to two fundamental exact sequences: the potential exact sequence and copotential exact sequence. Note that the potential exact sequence was first defined by DiPasquale in Generalized splines and graphic arrangements [4, Lemma 2.2].
@@ 4 Applications
Since $R$ is a domain, every non-zero principal ideal is generated by a regular element. Let $\mathfrak{m}$ be a maximal ideal of $R$. Now by the stability of spline structure morphisms under flat base change [8, Main Theorem] we have that $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$. That is, $(R_{G})_{\mathfrak{m}}$ is isomorphic to a module of splines over the local ring $R_{\mathfrak{m}}$ that has edge-labeled graph $G_{\mathfrak{m}}$. By hypothesis $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ satisfies the conditions of Theorem 3.5, and is therefore free.
Furthermore, each $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ is the kernel of a potential short exact sequence, for which the middle term is free, and since each $R_{\mathfrak{m}}$ is Noetherian [7, Lemma 10.31.1] we have that $(R_{G})_{\mathfrak{m}}$ is finitely generated. Now for modules over Noetherian rings, finitely generated implies finitely presented [7, Lemma 10.31.4] and therefore each $(R_{G})_{\mathfrak{m}}\cong(R_{\mathfrak{m}})_{G_{\mathfrak{m}}}$ is also finitely presented.
So we have established that for each maximal ideal $\mathfrak{m}$, we have that $(R_{G})_{\mathfrak{m}}$ is free and of finite presentation. Now any $R$-module that is free and of finite presentation locally at each maximal ideal of $R$ is also finite projective [7, Lemma 10.78.2(5)], so it follows that $R_{G}$ is a finite projective $R$-module as claimed. ∎
@@
Let $R$, $G$, and $R_{G}$ be given as specified. Note that by hypothesis $R$ is the coordinate ring of a nonsingular surface, and is therefore a regular ring [7, Definition 28.9.1, Lemma 28.9.2(3)]. Since regular rings are Cohen–Macaulay [5, Example 2.2], we can see $R$ is Cohen–Macaulay. We also know that $R$ is Krull dimension $2$ at each maximal ideal by hypothesis. Furthermore, since $R$ is a regular ring, it has finite global dimension, which implies that $R_{G}$ has finite projective dimension [7, Lemma 10.110.8, Definition 10.109.2]. Finally, by the irreducibility hypothesis we know $R$ is an integral domain, and it is Noetherian since the quotient of a Noetherian ring is Noetherian. So the criteria of Theorem 4.1 are satisfied and the claim follows. ∎